Problem 3 (4 point) Determine the x-coordinate of the centroid of the shaded area. у II 1 62 - b! - L x a
Answer #1

The x coordinate centroid of shaded area is  \frac{a}{2} + \frac{4\sqrt{2}}{3\pi }

Let's see how is it